# Mocap Everyone Everywhere: Lightweight Motion Capture With Smartwatches and a Head-Mounted Camera

### v1

@@ Abstract
We present a lightweight and affordable motion capture method based on two smartwatches and a head-mounted camera. In contrast to the existing approaches that use six or more expert-level IMU devices, our approach is much more cost-effective and convenient. Our method can make wearable motion capture accessible to everyone, enabling 3D full-body motion capture in diverse environments. As a key idea to overcome the extreme sparsity and ambiguities of sensor inputs, we integrate 6D head poses obtained from the head-mounted cameras for motion estimation. To enable capture in expansive indoor and outdoor scenes, we propose an algorithm to track and update floor level changes to define head poses, coupled with a multi-stage Transformer-based regression module. We also introduce novel strategies leveraging visual cues of egocentric images to further enhance the motion capture quality while reducing ambiguities. We demonstrate the performance of our method on various challenging scenarios, including complex outdoor environments and everyday motions including object interactions and social interactions among multiple individuals.
{strip}
Figure 1: We present a lightweight and affordable motion capture method based on two smartwatches and a head-mounted camera. Please visit our project page for a full visualization of the captured motion sequences.
@@ 1 Introduction
In this paper, we present a motion capture method that utilizes IMU sensors of smartwatches on wrists and a head-mounted monocular camera. We overcome the sparsity of using only 2 sensors and the intrinsic ambiguities of IMU sensors (orientation, acceleration) by integrating head 6DoF poses obtained from the head-mounted camera into the motion estimation pipeline. In contrast to VR settings (e.g., HMDs) where head poses are given in a fixed world coordinate in a small indoor environment [18, 30, 63, 37, 13], it is not trivial to define head poses in expansive outdoor settings. Through an algorithm that automatically tracks and adjusts floor levels, our approach can robustly capture motion in “in-the-wild” scenarios in challenging locations with non-flat grounds such as hilly areas or stairs. Furthermore, we also present a motion optimization approach for enhancing capture quality in complex everyday motions, such as object interactions, by leveraging observed body part cues through the head-mounted camera. As our system can be conveniently equipped by multiple users, we also explore scenarios where the signals among the individuals are shared and the egocentric observations from one person can be used as sparse third-person views for the others, which can be additionally used for our motion optimization module.
@@ 3.1 Overview
We build a transformer-based regression model coupled with a floor-level update module to estimate the full-body human motion from the cues of estimated head pose and wrist IMU signals: $\mathbf{M}_{reg}=\mathcal{F}_{est}(\mathbf{H},\mathbf{W},\mathbf{S}_{IMU})$. While our initial full-body estimation $\mathbf{M}_{reg}$ is already compelling, there exist fundamental ambiguities due to the extreme sparsity of our input signals. To address this issue, we leverage the visual cues captured by the head-mounted view $\mathbf{I}$, where the hand position and the interactions with the environments are observed. While available only occasionally, we demonstrate the use of such visual cues can enhance the motion capture quality. Going one step further, we also consider the multi-people scenarios where each individual wears our lightweight system for motion capture. Assuming the visual signals among people can be shared, we can additionally leverage the occasional “third-person” views from other people for the target person’s motion capture. To this end, we build a motion optimizer module by optimizing the human motion in spatiotemporal manifold space [26] to fuse all of these available signals for more accurate motion estimation: $\mathbf{M}=\mathcal{F}_{opt}(\mathbf{M}_{reg},\mathbf{\Phi})$, where $\mathbf{M}_{reg}$ is the initial output from our motion regressor module and $\mathbf{\Phi}$ is the visual cues extracted from head-mounted camera or third-person views from other users’ cameras. For $\mathbf{\Phi}$, we mainly use the 2D keypoint cues estimated from images in egocentric scenarios and 3D pose estimation results in multi-person scenarios.
@@ 3.3 Motion Estimation
Figure 4: Network architecture of submodules $\mathcal{F}^{end}$ (left) and $\mathcal{F}^{body}$ (right).
@@ 4.1 Experimental Setup
Baselines. Since we are the first to estimate full body motion from an egocentric camera and two IMU sensors, there exist no directly competing baselines. Thus, here, we compare our estimation module $\mathcal{F}_{est}$ with previous SOTAs based on (1) 6 IMUs (full body) [67, 31], and (2) VR devices (upper body, 6 DoF) [18], which pursue the similar goals with more cues than ours through additional sensors or 6DoF hand measurements. In the context of wearable and outdoor motion capture, we consider IMU-based methods as our main competitors. However, for completeness, we additionally compare with the VR device-based methods due to their technical similarity to our approach of estimating motion from upper body sensors. Note that in this comparison we assume fixed flat floor planes since previous approaches cannot handle varying floor levels.
@@ 4.2 Comparison with IMU-Based Baselines
Results. As demonstrated in Table 1, despite only utilizing sensors on the upper body, our method shows comparable or better performances compared to 6 IMU-based setups. In the root-relative position errors (r.MPJPE) which compare pose estimation quality ignoring root drifts, our approach outperforms the baselines. Notably, although PIP [67] explicitly corrected pose with physics-based optimization, our module shows comparable performance without such explicit corrections. In other metrics where root drift errors are also considered (MPJPE, Root PE), our method shows much better performance with significant decreases in root drift compared to baselines (MPJPE decreased by 86.1% for PIP and 88.5% for TIP; Root PE by 88.4% and 90.4%, respectively). This is expected as our method directly leverages head 6DoF poses for estimation. Yet, these results still demonstrate that incorporating head pose directly in estimation can mitigate root drifting issues without any additional localization or correction as done in previous approaches.
@@
Dataset Method MPJPE MPJVE Root PE Jitter
AMASS AGRoL [18] 4.58 19.31 4.15 2.60
AGRoL* 6.53 24.99 5.33 3.49
Ours ($\mathcal{F}_{est}$) 5.87 19.11 4.55 1.53
HPS AGRoL [18] 31.47 259.36 25.76 40.83
AGRoL* 33.40 389.72 28.21 62.06
Ours ($\mathcal{F}_{est}$) 10.08 26.08 7.69 1.67
Table 2: Comparison with upper body VR-based methods on AMASS and HPS datasets. AGRoL with an asterisk is the modified version where wrist position are replaced with accelerations.
@@ 4.3 Comparison with VR-Based Baselines
Dataset. For quantitative comparison, we generate train/test/validation split from a subset of AMASS dataset (CMU [57], HDM05 [48], BMLMovi [19]). We first evaluate our motion estimation module and AGRoL on the testing split of AMASS dataset. Additionally, we also demonstrate the generalization ability of the modules by testing on the HPS [23] dataset. Regarding HPS dataset, we use the motion optimization result [23] as ground truth. We randomly select 10 sequences in 6 scenes for comparison.
Results. As demonstrated in Table 2, our method shows on-par performance with the original AGRoL, even though AGRoL leverages the provided position values of both hands. Compared with the modified AGRoL where wrist positions are replaced with accelerations, our motion estimator module $\mathcal{F}_{est}$ outperforms it in all of the metrics. Furthermore, our module $F_{est}$ shows a notable performance gap compared to AGRoL on the HPS dataset which includes human motions in large-scale scenes. This is expected as AGRoL focuses on full body motion estimation in a VR-optimized setting, not in large-scale scenes.
@@ 4.4 Ablation Studies
Setup. For quantitative evaluation, we capture our own dataset by wearing XSens MVN Link [10] and our lightweight setups together, where we use XSens as ground truth data. However, as an IMU-based method XSens also suffers from root drift issues [23], and thus we ignore these factors during the evaluations, by considering errors in XY aligned space (in evaluating floor update algorithm) or capturing motion with relatively small root translations (in evaluating $F_{opt}$). For ablation on our floor level update algorithm, we capture a sequence of walking downstairs (height changes from 0m to -4.21m). For ablation on optimization $F_{opt}$, we capture 3 sequences involving hand-based human-scene interactions. (e.g., making coffee, taking snacks from the shelf) In the scenarios, the right hand is detected on the egocentric video and is targeted for optimization.
@@
Table 3: Quantitative comparison with $\mathcal{F}_{est}$ without floor level update and with floor level update.
@@
Ablation on Motion Estimation. We both quantitatively and qualitatively demonstrate the contribution of the floor level updating in scenarios where the floor changes over time (e.g., walking down the stairs). As seen in Table 3 and Fig. 5, without floor update (left) the head height is misleading and consequently, the module generates inaccurate poses. However, by accordingly updating the floor levels (right) the head height is adjusted so the network accurately estimates the motion of walking down the stairs.
Ablation on Motion Optimization. We quantitatively and qualitatively demonstrate the contribution of $\mathcal{F}_{opt}$ by comparing the initial motion output $\mathbf{M}_{reg}$ and optimized final motion output $\mathbf{M}$ to ground-truth data. As shown in Table 4, right arm PE and both arm PE both decreased in the final motion output $\mathbf{M}$ compared to the initial regression-based output $\mathbf{M}_{reg}$. Example results are shown in Fig. 6.
We also demonstrate motion optimization results from visual cues in multi-person scenarios. As seen in Fig. 7, the egocentric image of person B serves as a “third-person view” of person A. Using this information can resolve the fundamental ambiguities of the legs that stem from upper-body sensor setups. Although the legs of $\mathcal{F}_{est}$ (red) in Fig. 7 (a) are plausible, $\mathcal{F}_{opt}$ (blue) further captures the slight nuances of leg movements by leveraging visual cues $\bm{\tau}_{t}^{\overrightarrow{BA}}$.
@@ 5 Discussion
We present a novel, easy-to-use, and affordable motion capture system with smartwatches on wrists and a head-mounted camera. We overcome the sparsity and the ambiguity of sensor input by integrating head poses. By tracking and updating floor levels to define head poses coupled with a multi-stage Transformer-based estimation module our method can robustly capture motion in challenging locations with non-flat grounds. Furthermore, we also present a motion optimization approach by using visual cues through the head-mounted camera. As can be conveniently equipped by multiple users, we also explore scenarios where the signals among the individuals are shared.
@@ Acknowledgements.
This work was supported by SNU Creative-Pioneering Researchers Program and NRF grant funded by the Korea government (MSIT) (No. 2022R1A2C2092724 and No. RS-2023-00218601). H. Joo is the corresponding author.
@@ Appendix A Supplementary Video
The supplementary video shows the real-world demonstrations of our motion capture system with two smartwatches and a head-mounted camera. The real-world demonstrations include various scenarios including expansive outdoor scenes, everyday motions and interactions (e.g., making coffee), motions with dynamic movements, and social interactions among multiple people. We also visually demonstrate floor level changes in areas with drastic floor level changes such as walking down the stairs.
@@ B.1 Camera-Derived Head Poses
In the main paper, we have demonstrated the strength of our motion estimation module $\mathcal{F}_{est}$ against baselines, where we test with the head pose generated by available motion capture data given that paired egocentric videos are not available in the existing real IMU dataset. As an extension of this experiment, we further perform a similar quantitative evaluation with egocentric videos using the dataset in [66], where the motions in the TotalCapture [56] dataset are paired with egocentric videos taken from virtual head-mounted cameras in synthetic scenes. This additional experiment provides further evidence of the strength of our method in a closer scenario with the practical setup, accounting for potential noise introduced during the head pose estimation process in monocular SLAM.
We randomly select 30 sequences in total (about 60000 frames), 15 for each scene. We follow the procedure in Sec 3.2 to obtain head poses from the egocentric videos using DROID-SLAM [53]. As the dataset does not undergo the alignment procedure in our system beforehand (Supp. Sec. C), the camera poses obtained from SLAM are aligned to world coordinates using head poses derived from the dataset.
@@
In Table 6, we compare the motion estimation outputs with the camera-driven head poses, denoted as $\mathbf{H}^{C}+\mathcal{F}_{est}$, with the results with the methods with 6 IMU sensors, PIP [67] and TIP [31], and our method $\mathcal{F}_{est}$ with head poses from the dataset. Interestingly, our results with camera-driven head pose from the raw egocentric data mostly outperform the competing methods which use the 6 IMU sensors, even though our setup with an egocentric camera and fewer sensors (2 IMUs on the wrists) is much more challenging compared to the previous methods. The discrepancy in the results of the root-included metrics between $\mathbf{H}^{C}+\mathcal{F}_{est}$ and $\mathcal{F}_{est}$ can be considered due to the camera position errors, as depicted in Table 5.

### v2

@@ Abstract
We present a lightweight and affordable motion capture method based on two smartwatches and a head-mounted camera. In contrast to the existing approaches that use six or more expert-level IMU devices, our approach is much more cost-effective and convenient. Our method can make wearable motion capture accessible to everyone everywhere, enabling 3D full-body motion capture in diverse environments. As a key idea to overcome the extreme sparsity and ambiguities of sensor inputs with different modalities, we integrate 6D head poses obtained from the head-mounted cameras for motion estimation. To enable capture in expansive indoor and outdoor scenes, we propose an algorithm to track and update floor level changes to define head poses, coupled with a multi-stage Transformer-based regression module. We also introduce novel strategies leveraging visual cues of egocentric images to further enhance the motion capture quality while reducing ambiguities. We demonstrate the performance of our method on various challenging scenarios, including complex outdoor environments and everyday motions including object interactions and social interactions among multiple individuals. Project Page: https://jiyewise.github.io/projects/MocapEvery
{strip}[Uncaptioned image]
Figure 1: We present a lightweight and affordable motion capture method based on two smartwatches and a head-mounted camera.
@@ 1 Introduction
In this paper, we present a motion capture method that utilizes IMU sensors of smartwatches on wrists and a head-mounted monocular camera. We overcome the sparsity of using only 2 sensors and the intrinsic ambiguities of IMU sensors (orientation, acceleration) by integrating head 6DoF poses obtained from the head-mounted camera into the motion estimation pipeline. In contrast to VR settings (e.g., HMDs) where head poses are given in a fixed world coordinate in a small indoor environment [18, 30, 63, 37, 13], it is not trivial to define head poses in expansive outdoor settings. Through an algorithm that automatically tracks and adjusts floor levels, our approach can robustly capture motion in “in-the-wild" scenarios in challenging locations with non-flat grounds such as hilly areas or stairs. Furthermore, we also present a motion optimization approach for enhancing capture quality in complex everyday motions, such as object interactions, by leveraging observed body part cues through the head-mounted camera. As our system can be conveniently equipped by multiple users, we also explore scenarios where the signals among the individuals are shared and the egocentric observations from one person can be used as sparse third-person views for the others, which can be additionally used for our motion optimization module.
@@ 3.1 Overview
We build a transformer-based regression model coupled with a floor-level update module to estimate the full-body human motion from the cues of estimated head pose and wrist IMU signals: $\mathbf{M}_{reg}=\mathcal{F}_{est}(\mathbf{H},\mathbf{W},\mathbf{S}_{IMU})$. While our initial full-body estimation $\mathbf{M}_{reg}$ is already compelling, there exist fundamental ambiguities due to the extreme sparsity of our input signals. To address this issue, we leverage the visual cues captured by the head-mounted view $\mathbf{I}$, where the hand position and the interactions with the environments are observed. While available only occasionally, we demonstrate the use of such visual cues can enhance the motion capture quality. Going one step further, we also consider multi-people capture scenarios where each individual wears our lightweight system. Assuming the visual signals among people can be shared, we can additionally leverage the occasional “third-person” views from other people for the target person’s motion capture. To fuse all of these available signals for more accurate motion estimation, we build a motion optimizer module $\mathcal{F}_{opt}$ where $\mathbf{M}=\mathcal{F}_{opt}(\mathbf{M}_{reg},\mathbf{\Phi})$. Optimizing the human motion is done in a spatiotemporal manifold space [26], where $\mathbf{M}_{reg}$ is the initial output from our motion regressor module and $\mathbf{\Phi}$ is the visual cues extracted from head-mounted camera or third-person views from other users’ cameras. For $\mathbf{\Phi}$, we mainly use the 2D keypoint cues estimated from images in egocentric scenarios and 3D pose estimation results in multi-person scenarios.
@@ 3.3 Motion Estimation
Figure 4: Network architecture of submodules $\mathcal{F}^{end}$ (left) and $\mathcal{F}^{body}$ (right) of $\mathcal{F}_{est}$.
@@ 4.1 Experimental Setup
Baselines. Since we are the first to estimate full body motion from a head-mounted camera and two IMU sensors, no direct baselines exist. Thus, we compare our estimation module $\mathcal{F}_{est}$ with previous SOTAs based on (1) 6 IMUs (full body) [67, 31], and (2) VR devices (upper body, 6 DoF) [18], which pursue similar goals with more information than our setup through additional sensors or 6DoF hand measurements. In the context of wearable and outdoor motion capture, we consider IMU-based methods as our main competitors. However, for completeness, we additionally compare with the VR device-based methods due to their technical similarity to our approach of estimating motion from upper body sensors. Note that in this comparison we assume fixed flat floor planes since previous approaches cannot handle varying floor levels.
@@ 4.2 Comparison with IMU-Based Baselines
Results. As shown in Table 1, despite only utilizing sensors on the upper body, our method shows comparable or better performances compared to 6 IMU-based setups. In the root-relative position errors (r.MPJPE) which compare pose estimation quality ignoring root drifts, our approach outperforms the baselines. Notably, although PIP [67] explicitly correct pose with physics-based optimization, the performance of ours is comparable without such explicit corrections. In other metrics where root drift errors are also considered (MPJPE, Root PE), our method shows much better performance with significant decreases in root drift compared to baselines (MPJPE decreases by 87.7% for PIP and 89.8% for TIP; Root PE by 90.5% and 92.1%, respectively). For comparison with EgoLocate, results in supp. Table 2 shows our superior performance, especially in root-related position error terms (MPJPE, Root PE), despite the reduced number of sensors. These results demonstrate that incorporating head pose directly in estimation can improve performance and mitigate root drift issues without any additional localization or correction as done in previous approaches.
@@
Dataset Method MPJPE (r.) MPJVE Root PE Jitter
AMASS AGRoL [18] 4.58 (4.89) 19.31 4.15 2.60
AGRoL* 6.53 (6.18) 24.99 5.33 3.49
Ours ($\mathcal{F}_{est}$) 5.20 (4.95) 17.00 4.36 1.64
HPS AGRoL [18] 31.47 (27.64) 258.92 25.76 40.83
AGRoL* 33.46 (27.69) 389.83 28.22 62.06
Ours ($\mathcal{F}_{est}$) 8.65 (8.24) 21.42 6.97 1.84
Table 2: Comparison with upper body VR-based methods on AMASS and HPS datasets. AGRoL with an asterisk is the modified version where wrist position are replaced with accelerations. (r.) indicates root-relative position errors.
@@ 4.3 Comparison with VR-Based Baselines
Dataset. For quantitative comparison, we generate train/test/validation split from a subset of AMASS dataset (CMU [57], HDM05 [48], BMLMovi [19]). We first evaluate our motion estimation module and AGRoL on the testing split of the AMASS dataset. Additionally, we also demonstrate the generalization ability of the modules by testing on the HPS [23] dataset. Regarding HPS dataset, we use the motion optimization result [23] as ground truth. We randomly select 10 sequences in 6 scenes for comparison.
Results. As demonstrated in Table 2, our method shows on-par performance with the original AGRoL, even though in AGRoL the positions of both hands are provided. Compared with the modified AGRoL where wrist positions are replaced with accelerations, our module $\mathcal{F}_{est}$ outperforms it in all of the metrics. Furthermore, our module $\mathcal{F}_{est}$ shows a notable performance gap compared to AGRoL on the HPS dataset which includes motions in large-scale scenes. This is expected as AGRoL focuses on full body motion estimation in a VR-optimized setting, not in large-scale scenes.
@@ 4.4 Ablation Studies
Setup. For quantitative evaluation, we capture with our setup and with XSens MVN Link [10] together, where we use XSens as ground truth. More details are in supp. mat.
Ablation on Motion Estimation. We demonstrate the contribution of the floor-level updating algorithm by capturing a sequence of walking downstairs. As seen in Table 3 and Fig. 5, without floor update (left) the head height is misleading and consequently, the module generates inaccurate poses. However, by accordingly updating the floor levels (right) the head height is adjusted so the network accurately estimates the motion of walking down the stairs.
@@
Table 3: Quantitative comparison with and $\mathcal{F}_{est}$ without floor level update.
@@
Ablation on Motion Optimization. We first demonstrate the contribution of $\mathcal{F}_{opt}$ in hand-based everyday interaction scenarios. For quantitative comparison, we compare the initial motion output $\mathbf{M}_{reg}$ and optimized final output $\mathbf{M}$ to ground-truth. As shown in Table 4, right arm PE and both arm PE both decreased in the final motion output $\mathbf{M}$ compared to the initial regression-based output $\mathbf{M}_{reg}$. Example results are shown in Fig. 6. We also demonstrate motion optimization results from visual cues in multi-person scenarios. As seen in Fig. 7, the egocentric image of person B serves as a “third-person view" of person A. Using this information can resolve the fundamental ambiguities of the legs that stem from upper-body sensor setups. Although the legs of $\mathcal{F}_{est}$ (red) in Fig. 7 (a) are plausible, $\mathcal{F}_{opt}$ (blue) further captures the slight nuances of leg movements by leveraging visual cues $\bm{\tau}_{t}^{\overrightarrow{BA}}$.
@@ 5 Discussion
We present a novel, easy-to-use, and affordable motion capture system with smartwatches on wrists and a head-mounted camera. We overcome the sparsity and ambiguity of sensor inputs with different modalities by integrating head poses in the motion estimation pipeline. By tracking and updating floor levels to define head poses and incorporating into multi-stage Transformer-based estimation module our method can robustly capture motion in challenging locations with non-flat grounds. We further present a motion optimization approach by using visual cues through the camera. As multi-users can conveniently equip our setup, we also explore the scenarios when user signals are shared.
@@ Acknowledgements.
We thank Inhee Lee for supporting the system setup and thank Jiwon Song, Sumin Lee, and Hakjean Kim for assisting multi-person capture. This work was supported by SNU Creative-Pioneering Researchers Program, NRF grant funded by the Korea government (MSIT) (No.2022R1A2C2092724 and No. RS-2023-00218601), and IITP grant funded by the Korean government (MSIT) (No.2021-0-01343). H. Joo is the corresponding author.
@@ Appendix A Supplementary Video
The supplementary video in the project page shows the real-world demonstrations of our motion capture system with two smartwatches and a head-mounted camera. The real-world demonstrations include various scenarios including expansive outdoor scenes, everyday motions and interactions (e.g., making coffee), motions with dynamic movements, and social interactions among multiple people. We also visually demonstrate floor level changes in areas with drastic floor level changes such as walking down the stairs.
@@ B.1 Camera-Derived Head Poses
In the main paper, we have demonstrated the strength of our motion estimation module $\mathcal{F}_{est}$ against baselines, where we test with the head pose generated by available motion capture data given that paired egocentric videos are not available in the existing real IMU dataset. As an extension of this experiment, we further perform a similar quantitative evaluation with egocentric videos using the dataset in EgoLocate [66], where the motions in the TotalCapture [56] dataset are paired with egocentric videos taken from virtual head-mounted cameras in synthetic scenes. This additional experiment provides further evidence of the strength of our method in a closer scenario with the practical setup, accounting for potential noise introduced during the head pose estimation process in monocular SLAM. To demonstrate the reliability of the pre-processing stage of deriving head pose from egocentric video input, which happens prior to $\mathcal{F}_{est}$, we use the dataset in Egolocate [66] where the motions in the TotalCapture [56] dataset are paired with egocentric videos taken from virtual head-mounted cameras in synthetic scenes.
We randomly select 30 sequences in total (about 60000 frames), 15 for each scene. We follow the procedure in Sec. 3.2 to obtain head poses from the egocentric videos using DROID-SLAM [53]. As the dataset does not undergo the alignment procedure in our system beforehand (Supp. Sec. C), the camera poses obtained from SLAM are aligned to world coordinates using head poses derived from the dataset.
@@
In Table 6, we compare the motion estimation outputs with the camera-driven head poses, denoted as $\mathbf{H}^{C}+\mathcal{F}_{est}$, with the results with the methods with 6 IMU sensors, PIP [67] and TIP [31], and our method $\mathcal{F}_{est}$ with head poses from the dataset. Our results with camera-driven head pose from the raw egocentric data mostly outperform the competing methods which use the 6 IMU sensors, even though our setup with an egocentric camera and fewer sensors (2 IMUs on the wrists) is much more challenging compared to the previous methods.
We furthermore compare our results with EgoLocate [66], which leverages 6 IMU sensors on the full body for body pose estimation and an additional head-mounted camera for global translation localization. Interestingly, despite the reduced number of sensors, our results outperforms EgoLocate. Especially for root-related position error terms (MPJPE, Root PE) our method significantly outperforms. Such improvement, despite of reduced number of sensors, initially stems from the accuracy of camera pose estimation from DROID-SLAM [53] (Table 5), and by directly incorporating the 6DoF head pose cues from the cameras to the motion estimation process, unlike EgoLocate that employs a decoupled approach between motion estimation (IMU) and localization (camera).
The discrepancy in the results of the root-included metrics between $\mathbf{H}^{C}+\mathcal{F}_{est}$ and $\mathcal{F}_{est}$ can be considered due to the camera position errors, as depicted in Table 5.
@@
B.2 Ablation Studies on Motion Estimation
Although the motion estimation module $\mathcal{F}_{est}$ does not take the absolute position of wrists as input, our module $\mathcal{F}_{est}$ shows comparable or improved results compared to VR based baselines. (Table 2 of the main paper) To further clarify such result, we also conduct an ablation study on the motion estimation module to demonstrate the contribution of components in $\mathcal{F}_{est}$ for the performance. We first show the contribution of the multi-stage architecture (Sec. 3.3) where the initial module, $\mathcal{F}^{ee}$, determines end-effector positions before estimating the whole body motion. Moreover, we highlight the contribution of data representation in input which combines normalized and global coordinates. This combination of coordinates is advantageous as normalization standardizes actions and removes unwanted variations, while global information resolves ambiguities, like distinguishing between standing and sitting still. This dual-coordinate approach facilitates easier learning and enhances the network’s robustness in making inferences. Results in Table 7 validate the contribution of each component.